# Tony Lindeberg1    Zahra Babaiee2 and Peyman M. Kiasari2

### v1

@@ Abstract
This paper presents the results of analysing and modelling a set of 8 “master key filters”, which have been extracted by applying a clustering approach to the receptive fields learned in depthwise-separable deep networks based on the ConvNeXt architecture. For this purpose, we first compute spatial spread measures in terms of weighted mean values and weighted variances of the absolute values of the learned filters, which support the working hypotheses that: (i) the learned filters can be modelled by separable filtering operations over the spatial domain, and that (ii) the spatial offsets of the those learned filters that are non-centered are rather close to half a grid unit. Then, we model the clustered “master key filters” in terms of difference operators applied to a spatial smoothing operation in terms of the discrete analogue of the Gaussian kernel, and demonstrate that the resulting idealized models of the receptive fields show good qualitative similarity to the learned filters.
@@
Complementary experimental results then demonstrate the idealized models of receptive fields have good predictive properties for replacing the learned filters by idealized filters in depthwise-separable deep networks, thus showing that the learned filters in depthwise-separable deep networks can be well approximated by discrete scale-space filters.
Keywords:
receptive field deep learning discrete continuous Gaussian kernel Gaussian derivative depthwise-separable networks scale space
@@ 1 Introduction
In the area of deep learning, the problem of choosing the receptive fields is on the other hand addressed from a pure learning perspective, where the resulting receptive fields then arise as the result of optimizing some suitably selected loss function for a specific deep learning architecture. Hybrid approaches to this problem have also been formulated, where the use of Gaussian derivative operators have been extended to higher layers in deep networks by Jacobsen et al. (2016), Lindeberg (2021b, 2022), Pintea et al. (2021), Sangalli et al. (2022), Penaud-Polge et al. (2022), Gavilima-Pilataxi and Ibarra-Fiallo (2023) and Perzanowski and Lindeberg (2025). The use of Gaussian derivative operators in the higher layers of such hybrid networks is, however, not as theoretically motivated as the use of Gaussian derivatives in the first layer, formally based on axiomatically derived scale-space theory.
Filter 1 Filter 2 Filter 3 Filter 4
Refer to caption Refer to caption Refer to caption Refer to caption
Filter 5 Filter 6 Filter 7 Filter 8
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 1: The original set of 8 “master key filters”, obtained by applying a clustering technique to the receptive fields learned from depthwise-separable deep networks, as extracted by Babaiee et al. (2025a), “through greedy search on the ConvNeXt v2 Tiny model” developed by Liu et al. (2022).
Filter 1 Filter 2 Filter 3 Filter 4
Refer to caption Refer to caption Refer to caption Refer to caption
Filter 5 Filter 6 Filter 7 Filter 8
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 2: Alternative visualization of the original set of 8 “master key filters”, extracted by Babaiee et al. (2025a), while here complemented with a normalization the filters, (i) by rescaling Filters 1–6 to give the same response to matching first-order discrete monomials under discrete convolution as for convolution of the corresponding continuous monomials with corresponding continuous Gaussian derivative operators, according to (63)–(68), and (ii) by adding different constants to Filters 7–8, according to (69)–(73), to minimize the variance-based spatial spread measure of the filters, and finally visualizing the data on a blue-red colour scale with the value 0 corresponding to white, to better reveal the polarities of the filter values.
Idealized 1(A) Idealized 2(A) Idealized 3(A) Idealized 4(A)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(A) Idealized 6(A) Idealized 7(C2) Idealized 8(A)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method A in Section 3.1, based on direct transfer of scale values from the variances for a continuous Gaussian derivative model. (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{x,1},\sigma_{y,1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{x,2},\sigma_{y,2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{x,3},\sigma_{y,3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{x,4},\sigma_{y,4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{x,5},\sigma_{y,5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{x,6},\sigma_{y,6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{x,8},\sigma_{y,8})$. (The numerical values of the scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ for $i\in\{1,2,3,4,5,6,8\}$ have been determined according to Table 7, and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for the corresponding “master key filters”.)
@@
More recently, Babaiee et al. (2024b, 2024a, 2025a) have investigated the properties of depthwise convolutional filters learned across all layers of depthwise-separable convolutional neural networks, as originally proposed by Howard et al. (2017), with particular focus on the ConvNeXt architectures by Liu et al. (2022). Through unsupervised clustering techniques applied to millions of trained filters, they discovered that depthwise convolutional kernels can be categorized into a few distinct classes resembling Gaussian functions and their derivatives. Remarkably, these filter patterns demonstrate generality regardless of layer depth, network architecture, or training data domain. Through systematic greedy search methods, they further demonstrated that this diverse set of learned filters can without major loss of performance be reduced to a smaller set of just 8 so-called “master key filters”, see Figures 12 for two different ways of illustrating the same data.
@@
Notably, these receptive fields do specifically appear to be qualitatively very similar to receptive fields obtained from discrete scale-space theory, in terms of either (see Figure 3 for an illustration):
@@
In summary, the results to be presented will show that the learned filters in depthwise-separable deep networks based on the ConvNeXt architecture can be well modelled by discrete scale-space filters. By quantitative results regarding deep learning, we also show that replacement of the learned filter by idealized scale-space filters in depthwise-separable deep networks also lead to almost as good results as when using the fully learned filters. In these respects, the results presented in this paper generalize the previous solid theoretical foundation of using Gaussian derivative filters as the first layer of linear filters in a vision system to also using (here discrete approximations of) Gaussian derivative operators in the higher layers of deep networks.
Given that the ConvNeXt architecture has been demonstrated to constitute a modern architecture for deep networks with very good ability to compete with Transformer architectures, a main significance of this work is that it establishes a strong link between the filters learned in a reduced ConvNeXt architecture and the idealized filters predicted from theoretically based scale-space theory.
@@ 1.1 Structure of this paper
The presentation is organized as follows: Section 2 starts by describing a theoretical background for measuring characteristic properties of the learned filters, based on spatial spread measures defined from spatial means and spatial variances of the absolute values of the filter coefficients and the responses to lower-order polynomials of the learned filters. Section 3 then performs modelling of the 8 “master key filters” using four different types of main methods in terms of:
@@
Section 4 first interprets the results obtained in terms of the spaces spanned by the receptive field responses in terms of $N$-jet representations in terms of Gaussian derivatives as well as possible implications for the construction of more general Gaussian derivative networks, with the layers defined from linear combination of Gaussian derivative responses.
Section 5 then applies the fitted idealized discrete scale-space filters as the filtering primitives in depthwise-separable networks applied to the ImageNet dataset, and shows that (i) the idealized filters based on Method B have the predictive properties out of the four main classes of filter modelling methods, and (ii) using the idealized models of the 8 “master key filters” as the computational primitives in depthwise-separable deep networks lead to almost as good results as using the original learned filters. Finally, Section 6 concludes with a summary and discussion of some of the main results.

### v2

@@ Abstract
This paper presents the results of analysing and modelling a set of 8 “master key filters”, which have been extracted by applying a clustering approach to the receptive fields learned in depthwise-separable deep networks based on the ConvNeXt architecture.
For this purpose, we first compute spatial spread measures in terms of weighted mean values and weighted variances of the absolute values of the learned filters, which support the working hypotheses that: (i) the learned filters can be modelled by separable filtering operations over the spatial domain, and that (ii) the spatial offsets of the those learned filters that are non-centered are rather close to half a grid unit. Then, we model the clustered “master key filters” in terms of difference operators applied to a spatial smoothing operation in terms of the discrete analogue of the Gaussian kernel, and demonstrate that the resulting idealized models of the receptive fields show good qualitative similarity to the learned filters.
@@
Complementary experimental results then demonstrate that the idealized models of receptive fields have very good predictive properties for replacing the learned filters by idealized filters in depthwise-separable deep networks, thus showing that the learned filters in depthwise-separable deep networks can be well approximated by discrete scale-space filters. Notably, we show that, for a reduced version of the ConvNeXt architecture, using a set of only 8 discrete scale-space filters leads to almost as good accuracy as for the receptive fields trained from scratch on the ImageNet dataset.
@@ 1 Introduction
In the area of deep learning, the problem of choosing the receptive fields is on the other hand addressed from a pure learning perspective, where the resulting receptive fields then arise as the result of optimizing some suitably selected loss function for a specific deep learning architecture. Hybrid approaches to this problem have also been formulated, where the use of Gaussian derivative operators has been extended to higher layers in deep networks by Jacobsen et al. (2016), Lindeberg (2021b, 2022), Pintea et al. (2021), Sangalli et al. (2022), Penaud-Polge et al. (2022), Smets et al. (2023), Gavilima-Pilataxi and Ibarra-Fiallo (2023) and Perzanowski and Lindeberg (2025). The use of Gaussian derivative operators in the higher layers of such hybrid networks has, however, not previously been theoretically motivated by necessity, as the use of Gaussian derivatives in the first layer of a vision system has been formally established, based on the axiomatically derived scale-space theory by Iijima (1962), Koenderink (1984), Koenderink and van Doorn (1992), Lindeberg (1996, 2011, 2021a) and Weickert et al. (1999).
Filter 1 Filter 2 Filter 3 Filter 4 Filter 5 Filter 6 Filter 7 Filter 8
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 1: The original set of 8 “master key filters”, obtained by applying a clustering technique to the receptive fields learned from depthwise-separable deep networks, as extracted by Babaiee et al. (2025a), “through greedy search on the ConvNeXt V2 Tiny model” developed by Woo et al. (2023), in turn based on the regular ConvNeXt model developed by Liu et al. (2022). (Horizontal axes: horizontal filter indices $m\in[-3,3]$. Vertical axes: vertical filter indices $n\in[-3,3]$.)
Filter 1 Filter 2 Filter 3 Filter 4 Filter 5 Filter 6 Filter 7 Filter 8
Orig Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
B Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
C1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
C2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
D1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
D2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 2: Visualizations of the learned filters with the corresponding results of fitting idealized models of these filters using scale-space operations: (top row) Alternative visualization of the original set of 8 “master key filters”, extracted by Babaiee et al. (2025a), while here complemented with a normalization the filters, (i) by rescaling Filters 1–6 to give the same response to matching first-order discrete monomials under discrete convolution as for convolution of the corresponding continuous monomials with corresponding continuous Gaussian derivative operators, according to (63)–(68), and (ii) by adding different constants to Filters 7–8, according to (69)–(73), to minimize the variance-based spatial spread measure of the filters, and finally visualizing the data on a blue-red colour scale with the value 0 corresponding to white, to better reveal the polarities of the filter values. (rows 2-7) Idealized scale-space models of the filters, as computed with the different types of modelling approaches proposed in this paper: Method A in Section 4.1, based on direct transfer of scale values from the variances for a continuous Gaussian derivative model, Method B in Section 4.2, based on requiring the horizontal vs. the vertical discrete weighted variance-based spatial spread measures for the idealized receptive field models to be equal to the weighted discrete spatial spread measures for the corresponding learned filters, Method C1 in Section 4.3, based on minimizing the discrete $l_{1}$-norm between the idealized receptive field models and the normalized versions of the learned filters and using different values of the scale parameters in the horizontal and the vertical directions, Method C2 in Section 4.4, based on minimizing the discrete $l_{1}$-norm between the idealized receptive field models and the normalized versions of the learned filters and using the same values of the scale parameters in the horizontal and the vertical directions, Method D1 in Section 4.5, based on minimizing the discrete $l_{2}$-norm between the idealized receptive field models and the normalized versions of the learned filters and using different values of the scale parameters in the horizontal and the vertical directions, and Method D2 in Section 4.6, based on minimizing the discrete $l_{2}$-norm between the idealized receptive field models and the normalized versions of the learned filters and using the same values of the scale parameters in the horizontal and the vertical directions. (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for the corresponding “master key filters”.) (Horizontal axes: horizontal filter indices $m\in[-3,3]$. Vertical axes: vertical filter indices $n\in[-3,3]$.)
@@
More recently, Babaiee et al. (2024b, 2024a, 2025a) have investigated the properties of depthwise convolutional filters learned across all the layers of depthwise-separable convolutional neural networks, as originally proposed by Chollet (2017) and Howard et al. (2017), with particular focus on the ConvNeXt architectures by Liu et al. (2022) and Woo et al. (2023). Through unsupervised clustering techniques applied to millions of trained filters, they discovered that the depthwise convolutional kernels can be categorized into a few distinct classes, resembling Gaussian functions and their derivatives. Remarkably, these filter patterns demonstrate generality regardless of the layer depth, the network architecture, or the training data domain. Through systematic greedy search methods, they further demonstrated that this diverse set of learned filters can, without major loss of performance, be reduced to a smaller set of just 8 so-called “master key filters”, see Figure 1 and the first row in Figure 2 for two different ways of visualizing the same data. [^1] Concerning the visualizations of these filters, let us remark that that the actual scaling or the base levels of these filters is not regarded as essential for the functionality of the resulting deep networks in this study, since (i) the scaling of the filters can both be propagated across the layers in the network hierarchy with maintained essential functionality regarding the computations and can also be compensated for by rescalings of the weights that combine the output from the different filters, and (ii) the base levels of the filters can be compensated for with the bias terms in the linear combination of filter outputs before the non-linearity in each computational unit. Instead, for our approach, the spatial variations of the filters are regarded as providing the essence of the underlying computations, since these filter shapes determine how the local image information is to be combined between different spatial points in the image domain.
@@
Notably, these receptive fields do specifically appear to be qualitatively very similar to receptive fields obtained from discrete scale-space theory, in terms of either (see rows 2–7 in Figure 2 for illustrations):
@@
In summary, the results to be presented will show that the learned filters in depthwise-separable deep networks based on the ConvNeXt architecture can be well modelled by discrete scale-space filters. By quantitative results regarding deep learning, we also show that replacement of the learned filter by idealized scale-space filters in depthwise-separable deep networks also lead to almost as good results as when using the fully learned filters. In these respects, the results presented in this paper generalize the previous solid theoretical foundation of using Gaussian derivative filters as the first layer of linear filters in a vision system into a combined theoretical and experimentally based foundation to also using (here discrete approximations of) Gaussian derivative operators in the higher layers of deep networks.
Given that the ConvNeXt architecture has been demonstrated to constitute a modern state-of-the-art architecture for deep networks with very good ability to compete with Transformer architectures, a main significance of this work is that it establishes a strong link between the filters learned in a reduced ConvNeXt architecture and the idealized filters predicted from theoretically based scale-space theory.
1.1 Contributions and novelty
In brief, this paper comprises the following theoretical, conceptual and experimental contributions:
  • Extensions of the discrete scale-space theory to using non-centered discrete filter models.
  • A methodology for characterizing the properties of both idealized receptive field models and learned filters based on scale-space theory.
  • Methods for estimating the scale parameters in idealized receptive field models from learned filters based on matching of either continuous or discrete mean-based and variance-based spatial spread measures.
  • The formulation of weighted spatial spread measures, to reduce the risk that contributions from spurious values in the background may either lead to overestimates (biases) in the scale parameters.
  • The set of 8 idealized receptive field models obtained by model fitting for the 8 “master key filters”.
  • The proposal of replacing the learned filters in the ConvNeXt V2 architecture with idealized receptive fields based on scale-space theory, and evaluating the accuracy obtained when using different types of model fitting for determining the scale parameters in the idealized models of the receptive fields.
  • The result that model fitting based on discrete weighted spatial spread measures leads to significantly better predictive properties regarding the performance of using idealized receptive fields in a deep network architecture, compared to either using weighted continuous spatial spread measures or more traditional $l_{1}$- or $l_{2}$-based model fitting.
  • The result that using idealized receptive fields obtained from scale-space theory leads to almost as good accuracy as when using the originally learned filters in the ConvNeXt V2 Tiny architecture.
  • The result that only a marginal increase in the accuracy is obtained by replacing the values of the scale parameters obtained applying the best Method B to the 8 “master key filters” with the values of the scale parameters obtained by data-driven learning of the scale parameters using backpropagation.
  • A theoretically based interpretation of the spaces spanned by different sets of discrete derivative approximation operators applied to discrete Gaussian smoothing filters.
  • A set of predictions regarding the choices of the sets of filters to be used in further work regarding deep networks based on Gaussian derivatives or other closely related scale-space filters.
In these ways, the results provide additional support for using idealized scale-space filters as the receptive field primitives in deep networks, thereby extending (i) the previous axiomatic necessity results of using scale-space filters in the first layer in the visual hierarchy, as well as (ii) the previous sufficiency results of using Gaussian derivative kernels in higher layers of deep networks, towards (iii) showing that the receptive fields learned in all the layers of a modern state-of-the-art architecture for deep learning can be very well approximated by idealized scale-space filters.
@@ 1.2 Structure of this paper
The presentation is organized as follows: Section 2 provides a conceptual background regarding the ConvNeXt architecture for deep learning and the procedure for extracting the set of 8 “master key filters” from the filters learned from this architecture. Section 3 then starts by describing a theoretical background for measuring characteristic properties of the learned filters, based on spatial spread measures, defined from spatial means and spatial variances of the absolute values of the filter coefficients, as well as the responses to lower-order polynomials of the learned filters. Section 4 continues by performing modelling of the 8 “master key filters” using four different types of main methods in terms of:
@@
Section 5 then interprets the results obtained in this way in terms of the spaces spanned by the receptive field responses in terms of $N$-jet representations based on Gaussian derivatives. This section also describes implications and predictions from these results regarding the construction of more general Gaussian derivative networks, with the layers defined from linear combinations of Gaussian derivative responses.
Section 6 thereafter applies the fitted idealized discrete scale-space filters as the filtering primitives in depthwise-separable networks based on the ConvNeXt V2 Tiny architecture applied to the ImageNet dataset, and shows that:
  • the idealized filters based on Method B have the best predictive properties out of the four main classes of filter modelling methods, and
  • using the idealized models of the 8 “master key filters” as the computational primitives in depthwise-separable deep networks leads to almost as good results as if using the originally learned filters.
Finally, Section 7 concludes with a summary and discussion of some of the main results.
2 Background on depthwise-separable CNNs and the ConvNeXt architecture
2.1 Convolutional neural networks and depthwise-separable convolutions
Convolutional Neural Networks (CNNs) constitute a class of deep learning architectures in which spatial filtering operations are performed by learned convolutional kernels applied across multiple layers (LeCun et al. 1998). In a standard convolutional layer, each output channel is obtained by convolving all input channels with a kernel tensor of size $k\times k\times C_{\text{in}}$, where $k$ denotes the spatial kernel size and $C_{\text{in}}$ the number of input channels. Thus, spatial filtering and cross-channel mixing are performed jointly.
Depthwise-separable convolutions decouple these two operations into a spatial and a channel-mixing stage (Chollet 2017, Howard et al. 2017). Specifically:
  • Depthwise convolution: A spatial convolution with a $k\times k$ kernel is applied independently to each input channel. If the input has $C$ channels, this stage uses $C$ spatial kernels and does not mix information across channels.
  • Pointwise convolution: A so-called $1\times 1$ convolution is subsequently applied to linearly combine information across channels.
If both the input and output have $C$ channels, a standard convolution requires $k^{2}\,C^{2}$ parameters, whereas a depthwise-separable convolution requires $k^{2}\ C+C^{2}$ parameters. For moderate $k$ and large $C$, this leads to substantial computational savings.
From a structural viewpoint, the depthwise stage performs the entire spatial filtering operation. The subsequent pointwise convolution operates purely over the channel dimension.
2.2 The ConvNeXt architecture
ConvNeXt is a modern convolutional architecture designed to achieve competitive performance with Transformer-based vision models, while retaining a fully convolutional structure (Liu et al. 2022). It can be interpreted as a modernization of classical residual networks (He et al. 2016), incorporating architectural refinements inspired by Vision Transformers (Dosovitskiy et al. 2021).
A ConvNeXt block typically consists of:
  1. 1.
    a depthwise convolution (often with kernel size $7\times 7$),
  2. 2.
    a layer normalization (LN [^2] Instead of normalizing across the batch, like standard CNNs, the LayerNorm (LN) operation in the ConvNeXt V2 architecture normalizes across the channels for each individual pixel (or spatial location), helping to stabilize the training and mimicking the behavior of Vision Transformers.),
  3. 3.
    a pointwise ($1\times 1$) convolution expanding the channel dimension,
  4. 4.
    a nonlinearity (e.g., GELU [^3] The Gaussian Error Linear Unit (GELU) is a smoother non-monotonic replacement of the ReLU function, $\operatorname{GELU}(x)=\frac{x}{2}\left(1+\operatorname{erf}(\tfrac{x}{\sqrt{2}})\right)$, found to enhance the training in deep networks.),
  5. 5.
    a global response normalization (GRN [^4] Introduced in ConvNeXt V2, the global response normalization (GRN) layer enhances the feature diversity, by normalizing the activations of the channels based on their global relative importance, thereby preventing any single ”dead” or over-dominant channel from stalling the network.) layer, which enhances the feature diversity, by encouraging competition between the channels (Woo et al. 2023),
  6. 6.
    a second pointwise convolution, projecting back to the original channel dimension, and
  7. 7.
    a residual connection.
Crucially, the depthwise convolution is the only operation within the block that performs spatial filtering. All the subsequent operations act pointwise across spatial locations. Therefore, the spatial receptive field structure learned by ConvNeXt is entirely determined by its depthwise kernels.
Figure 3: Architectural overview of the ConvNeXt V2 Tiny network. Here, the abbreviation “LN” denotes “Layer Normalization”, while “GRN” denotes “Global Response Normalization”.
The global architecture of the ConvNeXt V2 network, that we will build our work upon, is organized into four hierarchical stages with a block distribution of $(3,3,9,3)$ for the Tiny variant, as illustrated in Figure 3. Starting from an initial patching layer with $C=96$ channels, with convolutional kernels and a stride size of $4$, the channel dimensions double at each stage ($96,192,384,768$), using strided convolutions for downsampling.
2.3 The Master Key Filters Hypothesis
Research shows that the trained filters in depthwise-separable CNNs consistently converge into a few identifiable clusters across all the layers (Babaiee et al. 2024b). Through unsupervised clustering of millions of filters, these patterns were found to resemble Gaussian functions and derivatives of Gaussian kernels. This discovery led to the Master Key Filters Hypothesis, which proposes that a general set of universal filters exists for visual data and that depthwise filters in depthwise-separable CNNs naturally tend toward these ”master keys” regardless of the specific dataset, architecture, or layer depth (Babaiee et al. 2025a).
To find the minimal set of the master key filters, an autoencoder was used to compress the trained $7\times 7$ depthwise kernels into a single-dimensional hidden code (Babaiee et al. 2025b). By sampling points from this latent 1D space, a spectrum of candidate filters was constructed. To find the minimal essential set, a greedy search was performed on the filters learned in the ConvNeXt V2 Tiny architecture, replacing its original filters with the closest linear transformations of these sampled candidates. The least important ones were removed iteratively, by measuring the change in accuracy, when each candidate was removed from the list. This process identified a distinct ”elbow point” on the accuracy plot at precisely 8 filters, where the accuracy remained stable up to that point despite a massive reduction in unique weights. These 8 essential filters, shown in Figure 1, can effectively replace thousands of trained parameters, while maintaining comparable performance on ImageNet.
2.4 Tangible contributions to ConvNeXt networks
In this work, we start from the given set of 8 “master key filters”, extracted using the procedure described above, and model these filters in terms of idealized models of receptive fields, as obtained from scale-space theory. By measuring and characterizing the spatial extents of these filters in terms of spatial spread measures, we show that the 8 “master key filters” can be well modelled by a set of 8 discrete scale-space filters.
By then replacing the originally learned filters in the ConvNeXt V2 Tiny architecture with these idealized models of spatial receptive fields, we show that it is possible to obtain almost as good results on the ImageNet dataset as if using the originally learned filters. This result thereby shows that the filters in the ConvNeXt V2 Tiny architecture can be very well approximated by discrete scale-space filters.